todd llc sells cars and trucks. the unit sales for trucks were $100,000 and the contribution margin was $200,000. the car unit sales were 120,000 and the contribution margin was $240,000. fixed costs are $220,000. how many trucks are needed to breakeven?
Answer #1

We know break even point BEP =Fixed cost/ Contribution margin per unit

Given

Trucks Contribution margin = 2,00,000

No of units = 1,00,000

Fixed cost = $220,000

Contribution margin per unit = Contribution margin / No of units = 2,00,000/1,00,000 = 2

Break even point BEP = 220000/2 = 1,10,000